The pitch of a screw jack is 0.5cm, when used to raise a load, the handle turns through a circle of 40cm. What is the mechanical advantage if the efficiency is 25%
1
Expert's answer
2020-04-15T10:40:22-0400
"VR=\frac{2\pi (40)}{0.5}=502.655"

"MA=(0.25)(502.655)=126"